OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair. (x'y) 2x+3y=12 3x+2y=13

OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. (3)2x + 3y = 12 and (-2) 3x + 2y = 13 6x+9y=36 and -6x-4y=-26 add together, x cancels out, 5y=10, y=2 plug back into one of 1st equations, 2x+3(2)=12-->2x=6, x=3 (x,y)=(3,2)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np haha